0 Required Intormation Problem 12-4A Partnershlp Income stetement of partners equlty, and closing entries LO P2 4% $31m.500, and $498.400, respectively They to La and S07000 io Barb, mierest siowances oco% en ther intial capital investments, and the balance shared as folows: 20% to Mo, 40%toLuond 40% to Bart. hiProblem 12-4A Port 2 O Type here to search
Answer #1

Answer

  • All working forms part of the answer
  • Requirement 2

Statement of Partner's Equity

Mo

Lu

Barb

Total

Initial partnership Investments

$               80,100

$             311,500

$          498,400

$            890,000

Net Income:

Salary Allowance

$               85,600

$               64,200

$            97,000

$            246,800

Interest (on capital) allowance

$                 8,010

$               31,150

$            49,840

$              89,000

Remaining Net Income (Loss)

$               37,540

$               75,080

$            75,080

$            187,700

Total Net Income

$            131,150

$             170,430

$          221,920

$            523,500

   TOTAL

$            211,250

$             481,930

$          720,320

$        1,413,500

Less: Drawings - Withdrawals

$               42,400

$               56,400

$            72,400

$            171,200

Ending Capital Investment balance

$            168,850

$             425,530

$          647,920

$        1,242,300

  • Working

Statement of Partner's Equity

Mo

Lu

Barb

Total

Initial partnership Investments

80100

311500

498400

=80100+311500+498400

Net Income:

Salary Allowance

85600

64200

97000

=85600+64200+97000

Interest (on capital) allowance

=80100*10%

=311500*10%

=498400*10%

=8010+31150+49840

Remaining Net Income (Loss)

=187700*20%

=187700*40%

=187700*40%

=523500-246800-89000

Total Net Income

Total

Total

Total

Total

TOTAL

=80100+131150

=311500+170430

=498400+221920

=211250+481930+720320

Less: Drawings - Withdrawals

42400

56400

72400

=42400+56400+72400

Ending Capital Investment balance

=211250-42400

=481930-56400

=720320-72400

=1413500-171200